OpenStudy (anonymous):

A 22-gauge copper wire has a radius of 0.645 mm, is 5.00 m long, and has a resistivity ρ = 1.72 × 10-8 Ωm. The maximum current for which this wire is rated is 0.920A. What is the maximum electric potential difference that may be applied to this wire without overloading it?

OpenStudy (anonymous):

@siddharth18

OpenStudy (anonymous):

Let max pot diff be V V = I * R R = (resistivity)(length of wire(in m)) / (Area of crossection) I = Current Area of crossection = pi * (radius)^2 Substitute the given quantities and solve

OpenStudy (anonymous):

Make sure you convert all quantities into SI units before substituting

OpenStudy (anonymous):

ho can i change from mm to m *10^3 or what ?

OpenStudy (anonymous):

\[1 mm = 10^{-3}m\]

OpenStudy (anonymous):

i am trying now to solve it can u give me the final answer to know if i solve correct or not

OpenStudy (anonymous):

oK. Ans is \[3.903 * 10^{-5}\]

OpenStudy (anonymous):

Volts

OpenStudy (anonymous):

are u sure i had to choose one of these answer 0.112 V 0.0605 V 0.561 V 0.242 V 0.342 V

OpenStudy (anonymous):

sorry. It is \[6.05*10^{-2}\]

OpenStudy (anonymous):

i.e option B. I forgot to square the radius :P

OpenStudy (anonymous):

:D thanx very much :)

OpenStudy (anonymous):

Welcome :)

OpenStudy (anonymous):

i will post another one now can u try to solve it :D

OpenStudy (anonymous):

A 1-m copper wire with a radius of 0.5 mm is stretched to a distance of 2 m. What is the fractional change in resistance ΔR/R as the wire is stretched?

OpenStudy (anonymous):

We know that(WKT) R = (resistivity)(length of wire(in m)) / (Area of crossection) ΔR/R= ((change in length)*(initial Area)) / (intial length)(change in area) Now volume of wire remains constant whle stretching. WKT vol = area* length (initial area)(initial length)=(final area)(final length) (Pi)(0.5 * 10^(-3) )^2 (1) = (final area)(2) final area = (pi) (0.25/2 * 10^(-6) ) = (pi) (0.125 * 10^(-6) ) Therefore ΔR/R= ((change in length)*(initial Area)) / (intial length)(change in area) = ( 1* (Pi)(0.25 * 10^(-6) ) / (1 * (pi) ( 0.5-0.125 ) * 10^ (-6) Now simplify. You will get ΔR/R= 1.33 I hope I have not made silly mistakes now :P